# Cosmological perturbations of a relativistic MOND theory

### v1

@@ IV.1.1 Gauge transformation to 1PN order
From the gauge transformation of the metric $g_{0i}$, under our metric ansatz in Eq. (21), we have $\xi^{(2)0}=\xi^{(2)0}(t)$ Hwang-Noh-Puetzfeld-2008 . The spatially constant $\xi^{(2)0}$ can be absorbed by a global redefinition of the time coordinate, and without losing generality we can set $\xi^{(2)0}\equiv 0$; this is necessary to maintain $V=U$, thus $\Psi=\Phi$ to 0PN order Hwang-Noh-Puetzfeld-2008 . Therefore, $\sigma$ cannot be used to fix the temporal gauge condition; in another words, to 0PN order, the PN approximation does not depend on the gauge Hwang-Noh-Puetzfeld-2008 . To 1PN order $\Sigma\equiv 0$ is a legitimate gauge condition which removes the gauge mode completely.
Still, we may set $\sigma=0$ as a physical condition on the $\tau$-fluid Bonetti-2015 . In this case, Eq. (31) implies $v_{i}=0$, thus comoving with the $\tau$-fluid. To 1PN order, we expect $\Sigma=0$ also implies comoving with the $\tau$-fluid, see $T^{0}_{i}$ in Eq. (54). In perturbation theory later, $\sigma$ and $\Sigma$ can be used as the temporal gauge condition which removes the gauge degree of freedom completely; $\sigma\equiv 0\equiv\Sigma$ corresponds to the comoving gauge of the $\tau$-fluid, see Sec. V.1.1. In our PN approximation, however, condition $\sigma=0$ to 0PN order should be considered as a physical condition on the $\tau$-field.
@@ IV.1.2 MOND regime with $\sigma=0$
$A={1\over a^{2}c^{4}}\Phi^{,i}\Phi_{,i}={1\over a^{2}c^{4}}|\nabla\Phi|^{2}={1\over a^{2}c^{4}}a_{M}^{2}x^{2}.$ (45)
In order to achieve the MOND and the proper Newtonian gravity in the low and high acceleration regimes, respectively, we demand the following. (i) In the MOND regime, $f(x)=x$ for $x\ll 1$ (MOND regime), we need ${\cal J}_{,A}=x-1=ac^{2}\sqrt{A}/a_{M}-1$, thus Blanchet-Skordis-2024
${\cal J}=-A+{2\over 3}{ac^{2}\over a_{M}}A^{3/2}+{\cal O}(x^{4}).$ (46)
(ii) In the Newtonian regime, we need $f(x)=1$ for $x\gg 1$. This implies ${\cal J}_{,A}$ should be negligibly small compared with unity in that limit Blanchet-Skordis-2024 . For a simple interpolating function $f(x)=x/(1+x)$, we have ${\cal J}_{,A}=-1/(1+x)$ and
${\cal J}=-2{a_{M}\over ac^{2}}\Big{[}\sqrt{A}-{a_{M}\over ac^{2}}\ln{\Big{(}1+{ac^{2}\over a_{M}}\sqrt{A}\Big{)}}\Big{]}.$ (47)
In galactic and smaller scales where MOND phenomena apply the cosmic expansion effect is negligible and we may set $a\equiv 1$. Setting $\tau=t$ for the background is consistent with this. In expanding cosmological background, as in Sec. III, $\tau$-field with $\tau(t)$ contributes to the background evolution through ${\cal K}$. In our PN approximation, as we consider $\tau=t$ for the background, the $\tau$-field does not contribute to the background. For $\tau=\tau(t)$ in the background, it is unclear whether a proper PN expansion is possible. Similarly, the requirement of a proper PN expansion constrains allowed functional form of ${\cal K}$.
The MOND regime naturally appears by choosing ${\cal J}$-function in Eq. (46). Although stationary condition is mentioned in Flanagan-2023 ; Blanchet-Skordis-2024 , we did not impose such a condition. We emphasize that setting $\sigma=0$ is a physical condition imposed on the $\tau$-field.
@@ IV.1.3 MOND constraint for $n=1$
To 1PN order, for simplicity we set $\sigma=0=\Sigma$. From Eq. (24), we have
@@
$U_{0}=-\Big{(}1+{1\over c^{2}}\Phi-{1\over 2c^{4}}\Phi^{2}\Big{)},\quad U_{i}=0,$
$U^{0}=1-{1\over c^{2}}\Phi+{3\over 2c^{4}}\Phi^{2},\quad U^{i}={1\over c^{3}}{1\over a}P^{i};$
$A_{0}=-{1\over c^{3}}{1\over a}P^{i}\Phi_{,i},\quad A_{i}=\Phi_{,i}\Big{(}1-{2\over c^{2}}\Phi\Big{)},$
$A^{0}=0,\quad A^{i}={1\over a^{2}}\Phi^{,i}\Big{[}1+{2\over c^{2}}(\Psi-\Phi)\Big{]},$
$A={1\over c^{4}}{1\over a^{2}}\Phi^{,i}\Phi_{,i}\Big{[}1+{2\over c^{2}}(\Psi-2\Phi)\Big{]}.$ (50)
@@
${\cal K}={1\over c^{4}}\nu^{2}\Phi^{2},\quad Q{\cal K}_{,Q}=-{2\over c^{2}}\nu^{2}\Phi\Big{[}1-{1\over c^{2}}\Big{(}{5\over 2}+{\cal K}_{3}\Big{)}\Phi\Big{]}.$ (51)
@@
$\hskip-14.22636pt\mu={1\over 8\pi G}\Big{\{}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}(c^{2}+v^{2}+2\Psi-2\Phi)+c^{4}{\cal J}+{2\over a^{2}}{\cal J}_{,A}(\Phi+\Psi)^{,i}\Phi_{,i}-2\nu^{2}\Phi\big{[}c^{2}+v^{2}-(2+{\cal K}_{3})\Phi\big{]}\Big{\}},$
$\hskip-14.22636ptp={1\over 8\pi G}{1\over 3}\Big{[}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}v^{2}-3c^{4}{\cal J}+{2\over a^{2}}{\cal J}_{,A}\Phi^{,i}\Phi_{,i}-\nu^{2}\Phi(2v^{2}-3\Phi)\Big{]},$
$\hskip-14.22636ptQ_{i}={1\over 8\pi G}\Big{\{}{2\over a^{2}}v_{i}({\cal J}_{,A}\Phi^{,j})_{,j}\Big{(}c^{2}+{3\over 2}v^{2}+2\Psi-2\Phi\Big{)}-{2\over a^{2}}{\cal J}_{,A}\big{[}\Phi_{,i}\Phi^{,j}v_{j}+v_{i}(\Phi+\Psi)^{,j}\Phi_{,j}\big{]}$
@@
$\hskip-14.22636pt\Pi_{ij}={1\over 8\pi G}\Big{[}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,k})_{,k}\Big{(}v_{i}v_{j}-{1\over 3}v^{2}\delta_{ij}\Big{)}+{2\over a^{2}}{\cal J}_{,A}\Big{(}\Phi_{,i}\Phi_{,j}-{1\over 3}\Phi^{,k}\Phi_{,k}\delta_{ij}\Big{)}-2\nu^{2}\Phi\Big{(}v_{i}v_{j}-{1\over 3}v^{2}\delta_{ij}\Big{)}\Big{]}.$ (52)
To 1PN order, we have $\mu=\varrho(c^{2}+\Pi)$ with $\varrho\Pi$ the internal energy density Chandrasekhar-1965 . The energy-frame condition $q_{i}\equiv 0$ for the $\tau$-fluid implies $v_{i}=0$, thus the velocity four-vector becomes normal, $u_{a}=n_{a}$. The fluid quantities become
$\mu={1\over 8\pi G}\Big{\{}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}(c^{2}+2\Psi-2\Phi)+c^{4}{\cal J}+{2\over a^{2}}{\cal J}_{,A}(\Phi+\Psi)^{,i}\Phi_{,i}-2\nu^{2}\Phi\big{[}c^{2}-(2+{\cal K}_{3})\Phi\big{]}\Big{\}},$
$p={1\over 8\pi G}{1\over 3}\Big{(}-3c^{4}{\cal J}+{2\over a^{2}}{\cal J}_{,A}\Phi^{,i}\Phi_{,i}+3\nu^{2}\Phi^{2}\Big{)},\quad\Pi_{ij}={1\over 8\pi G}{2\over a^{2}}{\cal J}_{,A}\Big{(}\Phi_{,i}\Phi_{,j}-{1\over 3}\Phi^{,k}\Phi_{,k}\delta_{ij}\Big{)}.$ (53)
@@
$T^{0}_{0}=-\mu,\quad T^{0}_{i}=0,\quad T^{i}_{0}=-{1\over c^{3}a}\mu P^{i},\quad T^{i}_{j}=p\delta^{i}_{j}+\Big{(}1+{2\over c^{2}}\Psi\Big{)}\Pi^{i}_{j}.$ (54)
We note that setting $q_{i}=0$ is a frame choice always allowed for any energy-momentum component. Under this choice, the flux or momentum of the component is encoded in the four-vector $u_{a}$. As the four-vector becomes normal with $n_{i}\equiv 0$, we have the physical momentum of the component vanish; this is reflected in $T^{0}_{i}=(\mu+p)u^{0}u_{i}=0$ to 1PN order. This implies that the $\tau$-fluid quantities are measured by an observer comoving with the $\tau$-fluid. This happened as we imposed $\sigma=0=\Sigma$ as a physical condition.
@@
Equation (60) is valid to 0PN order whereas the other two equations are valid to 1PN order. To 0PN order, Eq. (59) gives Eq. (38). To 0PN order, Eq. (60) is the same as Eq. (59), with $\Psi=\Phi$.
The fluid quantities in Eqs. (59)-(61) are the sum of other fluids besides the $\tau$-fluid; we may include the $\tau$-fluid to the collective fluid quantities as $\mu=\mu_{m}+\mu_{\tau}$, $p=p_{m}+p_{\tau}$, and $\varrho v_{i}=\varrho^{m}v^{m}_{i}+\varrho^{\tau}v^{\tau}_{i}$. For the $\tau$-fluid we have conservation equations in (55)-(56) or the equation of motion in Eq. (57). The fluid conservation equations to 1PN order are in Eqs. (B23) and (B24) of Hwang-Noh-axion-2023 . These are the complete set of equations valid to 1PN order without imposing the temporal gauge condition; for the choice of gauge conditions, see Sec. 6 in Hwang-Noh-Puetzfeld-2008 .

### v2

@@ IV.1.1 Gauge transformation to 1PN order
From the gauge transformation of the metric $g_{0i}$, under our metric ansatz in Eq. (21), we have $\xi^{(2)0}=\xi^{(2)0}(t)$ Hwang-Noh-Puetzfeld-2008 . The spatially constant $\xi^{(2)0}$ can be absorbed by a global redefinition of the time coordinate, and without losing generality we can set $\xi^{(2)0}\equiv 0$; this is necessary to maintain $V=U$, thus $\Psi=\Phi$ to 0PN order Hwang-Noh-Puetzfeld-2008 . Therefore, $\sigma$ cannot be used to fix the temporal gauge condition; in another words, to 0PN order, the PN approximation does not depend on the gauge which is natural as Newtonian gravity is free from the gauge issue Hwang-Noh-Puetzfeld-2008 . To 1PN order $\Sigma\equiv 0$ is a legitimate gauge condition which removes the gauge mode completely.
Still, we may set $\sigma=0$, but it is a physical condition on the $\tau$-fluid Bonetti-2015 ; it corresponds to setting $v_{i}=0$ which is a serious condition on the fluid behavior of the $\tau$-field; $v_{i}=0$ implies comoving with the $\tau$-fluid. The energy conservation equation in (36) leads to $\varrho\propto a^{-3}$ for the $\tau$-fluid. To 1PN order, we expect $\Sigma=0$ also implies comoving with the $\tau$-fluid, see $T^{0}_{i}$ in Eq. (61). In perturbation theory later, $\sigma$ and $\Sigma$ can be used as the temporal gauge condition which removes the gauge degree of freedom completely; $\sigma\equiv 0\equiv\Sigma$ corresponds to the comoving gauge of the $\tau$-fluid, see Sec. V.1.1. In our PN approximation, however, condition $\sigma=0$ to 0PN order should be considered as a physical condition on the $\tau$-field.
@@ IV.1.2 MOND regime with $\sigma=0$
$A={1\over c^{4}a^{2}}\Phi^{,i}\Phi_{,i}={1\over c^{4}a^{2}}|\nabla\Phi|^{2}={a_{M}^{2}\over c^{4}}x^{2}.$ (45)
In order to achieve the MOND and the proper Newtonian gravity in the low and high acceleration regimes, respectively, we demand the following.
(i) In the MOND regime, $f(x)=x$ for $x\ll 1$, we need ${\cal J}_{,A}=x-1=c^{2}\sqrt{A}/a_{M}-1$, thus Blanchet-Skordis-2024
${\cal J}=-A+{2\over 3}{c^{2}\over a_{M}}A^{3/2}+{\cal O}(x^{4}).$ (46)
(ii) In the Newtonian regime, we need $f(x)=1$ for $x\gg 1$. This implies ${\cal J}_{,A}$ should be negligibly small compared with unity in that limit Blanchet-Skordis-2024 .
For a simple interpolating function $f(x)=x/(1+x)$, we have
${\cal J}_{,A}=-{1\over 1+x},$ (47)
and
${\cal J}=-2{a_{M}\over c^{2}}\Big{[}\sqrt{A}-{a_{M}\over c^{2}}\ln{\Big{(}1+{c^{2}\over a_{M}}\sqrt{A}\Big{)}}\Big{]}.$ (48)
In galactic and smaller scales where MOND phenomena apply, the cosmic expansion effect is negligible and we may set $a\equiv 1$. In expanding cosmological background, as in Sec. III, $\tau$-field with $\tau(t)$ contributes to the background evolution through ${\cal K}$. In our PN approximation, as we consider $\tau=t$ for the background, the $\tau$-field does not contribute to the background. For $\tau=\tau(t)$ in the background, it is unclear whether a proper PN expansion is possible. The requirement of a proper PN expansion may constrain allowed functional form of ${\cal K}$.
The MOND regime naturally appears by choosing ${\cal J}$-function in Eq. (46). Still, we emphasize that setting $\sigma=0$ is a physical condition imposed on the $\tau$-field.
@@ IV.1.3 MOND constraint for $n=1$
IV.1.4 Faster growth in MOND regime
Here we show that the baryon density perturbation grows faster in the MOND regime compared with the Newtonian case. Equations (36) and (37) are valid for baryon with $p=0=\Pi_{ij}$. To the linear order, we have
$\dot{\delta}_{b}+{1\over a}\nabla\cdot{\bf v}_{b}=0,\quad{1\over a}(a{\bf v}_{b})^{\displaystyle{\cdot}}=-{1\over a}\nabla\Phi,$ (50)
thus
$\ddot{\delta}_{b}+2H\dot{\delta}_{b}={\Delta\over a^{2}}\Phi.$ (51)
Imposing $\sigma=0$, to the linear order, the MOND-modified Poisson’s equation in Eq. (43) gives
$(1+{\cal J}_{,A}){\Delta\over a^{2}}\Phi=4\pi G\delta\varrho_{b},$ (52)
where we ignored $\nu^{2}$-term from $n=1$ as it is excluded in Eq. (49). Combining these, we have
$\ddot{\delta}_{b}+2H\dot{\delta}_{b}={4\pi G\varrho_{b}\delta_{b}\over 1+{\cal J}_{,A}}=\Big{(}1+{1\over x}\Big{)}4\pi G\varrho_{b}\delta_{b},$ (53)
where in the second step we used the MOND function in Eq. (47) with $x\equiv{1\over a}|\nabla\Phi|/a_{M}$. In the MOND regime we have $x\ll 1$, and the gravity affecting density perturbation is stronger than in the Newtonian regime with $x\gg 1$.
For a pressureless background with $\Lambda=0$, in the MOND regime we have a solution with ${1\over a}\nabla\Phi={1\over a_{0}}\nabla\Phi_{0}$ and
${1\over a}|\nabla\Phi|={3\over 10}a_{M}\Omega_{b0},\quad\delta_{b}={{3\over 10}\Omega_{b0}{1\over a}\Delta\Phi\over 4\pi G\varrho_{b}a^{3}}a^{2}.$ (54)
Thus, $\delta_{b}\propto a^{2}$ grows faster than in the CDM case where $\delta_{b}\propto a$.
The faster growth of galaxies was previously noticed in the MOND paradigm Sanders-1998 ; McGaugh-2024 ; in the low-acceleration MOND regime, we have $|\vec{a}|=\sqrt{a_{M}|\vec{g}_{N}|}>|\vec{g}_{N}|$, thus the gravity is stronger than Newtonian gravity, and the structures can grow faster. Here, we showed the case using the BMS model and our result agrees with a heuristic derivation in Nusser-2002 .
@@ IV.1.4 Faster growth in MOND regime
To 1PN order, for simplicity we set $\sigma=0=\Sigma$; $\sigma=0$ is a physical condition we impose whereas $\Sigma=0$ is a temporal gauge condition we choose. From Eq. (24), we have
@@
$U_{0}=-\Big{(}1+{1\over c^{2}}\Phi-{1\over 2c^{4}}\Phi^{2}\Big{)},\quad U_{i}=0,\quad U^{0}=1-{1\over c^{2}}\Phi+{3\over 2c^{4}}\Phi^{2},\quad U^{i}={1\over c^{3}}{1\over a}P^{i};\quad A_{0}=-{1\over c^{3}}{1\over a}P^{i}\Phi_{,i},$
$A_{i}=\Phi_{,i}\Big{(}1-{2\over c^{2}}\Phi\Big{)},\quad A^{0}=0,\quad A^{i}={1\over a^{2}}\Phi^{,i}\Big{[}1+{2\over c^{2}}(\Psi-\Phi)\Big{]},\quad A={1\over c^{4}}{1\over a^{2}}\Phi^{,i}\Phi_{,i}\Big{[}1+{2\over c^{2}}(\Psi-2\Phi)\Big{]}.$ (56)
@@
${\cal K}={1\over c^{4}}\nu^{2}\Phi^{2}\Big{(}1-{3\over c^{2}}\Phi\Big{)},\quad Q{\cal K}_{,Q}=-{2\over c^{2}}\nu^{2}\Phi\Big{[}1-{1\over c^{2}}\Big{(}{5\over 2}+{\cal K}_{3}\Big{)}\Phi\Big{]}.$ (57)
@@
Using Eq. (4), Eq. (5) for the $\tau$-field gives
$T_{00}={c^{2}\over 8\pi G}\Big{[}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}\Big{(}1+{2\over c^{2}}\Psi\Big{)}+{2\over a^{2}c^{2}}{\cal J}_{,A}(\Phi+\Psi)_{,i}\Phi^{,i}-2\nu^{2}\Phi\Big{(}1-{1\over c^{2}}{\cal K}_{3}\Phi\Big{)}+{\cal J}c^{2}\Big{]},\quad T_{0i}=0,$
$T_{ij}={1\over 8\pi G}\Big{\{}a^{2}\Big{[}\nu^{2}\Phi^{2}\Big{(}1-{3\over c^{2}}\Phi-{2\over c^{2}}\Psi\Big{)}-{\cal J}c^{4}\Big{(}1-{2\over c^{2}}\Psi\Big{)}\Big{]}\delta_{ij}+2{\cal J}_{,A}\Phi_{,i}\Phi_{,j}\Big{(}1-{4\over c^{2}}\Phi\Big{)}\Big{\}}.$ (58)
@@
$\hskip-14.22636pt\mu={1\over 8\pi G}\Big{\{}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}(c^{2}+v^{2}-2\Phi+2\Psi)+{2\over a^{2}}{\cal J}_{,A}(\Phi+\Psi)^{,i}\Phi_{,i}-2\nu^{2}\Phi\big{[}c^{2}+v^{2}-(2+{\cal K}_{3})\Phi\big{]}+{\cal J}c^{4}\Big{\}},$
$\hskip-14.22636ptp={1\over 8\pi G}{1\over 3}\Big{\{}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}\Big{[}1+{1\over c^{2}}(v^{2}-2\Phi+4\Psi)\Big{]}v^{2}-2\nu^{2}\Phi\Big{[}1+{1\over c^{2}}\Big{(}v^{2}-{5\over 2}\Phi-{\cal K}_{3}\Phi+2\Psi\Big{)}\Big{]}v^{2}+3\nu^{2}\Phi^{2}\Big{(}1-{3\over c^{2}}\Phi\Big{)}$
$\qquad\hskip-14.22636pt+{2\over a^{2}}{\cal J}_{,A}\Phi^{,i}\Phi_{,i}\Big{[}1+{1\over c^{2}}(2\Psi-4\Phi)\Big{]}+{2\over a^{2}c^{2}}{\cal J}_{,A}(\Phi+\Psi)_{,i}\Phi^{,i}v^{2}+{2\over a^{2}c^{2}}{\cal J}_{,A}\Phi_{,i}v^{i}\Phi_{,j}v^{j}-3{\cal J}c^{4}\Big{\}},$
$\hskip-14.22636ptQ_{i}={1\over 8\pi G}\Big{\{}{2\over a^{2}}v_{i}({\cal J}_{,A}\Phi^{,j})_{,j}\Big{(}c^{2}+{3\over 2}v^{2}-2\Phi+2\Psi\Big{)}-{2\over a^{2}}{\cal J}_{,A}\big{[}\Phi_{,i}\Phi^{,j}v_{j}+v_{i}(\Phi+\Psi)^{,j}\Phi_{,j}\big{]}$
@@
$\hskip-14.22636pt\Pi_{ij}={1\over 8\pi G}\Big{\{}-\Big{(}v_{i}v_{j}-{1\over 3}\delta_{ij}v^{2}\Big{)}\Big{[}{2\over a^{2}}({\cal J}_{,A}\Phi^{,k})_{,k}\Big{(}1-{2\over c^{2}}\Phi+{2\over c^{2}}\Psi+{1\over c^{2}}v^{2}\Big{)}+2\nu^{2}\Phi\Big{(}1+{1\over c^{2}}v^{2}-{5\over 2c^{2}}\Phi-{1\over c^{2}}{\cal K}_{3}\Phi\Big{)}$
$\qquad\hskip-14.22636pt-{2\over a^{2}c^{2}}{\cal J}_{,A}(\Phi+\Psi)_{,k}\Phi^{,k}\Big{]}+{2\over a^{2}}{\cal J}_{,A}\Big{(}\Phi_{,i}\Phi_{,j}-{1\over 3}\delta_{ij}\Phi^{,k}\Phi_{,k}\Big{)}\Big{(}1-{4\over c^{2}}\Phi\Big{)}+{2\over a^{2}c^{2}}{\cal J}_{,A}\Phi_{,k}v^{k}\Big{(}v_{i}\Phi_{,j}+v_{j}\Phi_{,i}-{2\over 3}\delta_{ij}\Phi_{,\ell}v^{\ell}\Big{)}$
$\qquad\hskip-14.22636pt-{2\over 3c^{2}}v_{i}v_{j}\Big{[}{2\over a^{2}}({\cal J}_{,A}\Phi^{,k})_{,k}v^{2}+2\nu^{2}\Phi v^{2}+{1\over a^{2}}{\cal J}_{,A}\Phi^{,k}\Phi_{,k}\Big{]}+{2\over 3c^{2}a^{2}}\delta_{ij}{\cal J}_{,A}\Phi_{,k}v^{k}\Phi_{,\ell}v^{\ell}\Big{\}}.$ (59)
To 1PN order, we have $\mu=\varrho(c^{2}+\Pi)$ with $\varrho\Pi$ the internal energy density Chandrasekhar-1965 and $\Pi^{i}_{i}={1\over c^{2}}\Pi_{ij}v^{i}v^{j}$. The energy-frame condition $q_{i}\equiv 0$ for the $\tau$-fluid implies $v_{i}=0$, thus the velocity four-vector becomes normal, $u_{a}=n_{a}$. The fluid quantities become
$\mu={1\over 8\pi G}\Big{\{}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}(c^{2}-2\Phi+2\Psi)+{2\over a^{2}}{\cal J}_{,A}(\Phi+\Psi)^{,i}\Phi_{,i}-2\nu^{2}\Phi\big{[}c^{2}-(2+{\cal K}_{3})\Phi\big{]}+{\cal J}c^{4}\Big{\}},$
$p={1\over 8\pi G}\Big{[}\nu^{2}\Phi^{2}\Big{(}1-{3\over c^{2}}\Phi\Big{)}+{2\over 3a^{2}}{\cal J}_{,A}\Phi^{,i}\Phi_{,i}\Big{(}1-{4\over c^{2}}\Phi+{2\over c^{2}}\Psi\Big{)}-{\cal J}c^{4}\Big{]},$
$\Pi_{ij}={1\over 8\pi G}{2\over a^{2}}{\cal J}_{,A}\Big{(}\Phi_{,i}\Phi_{,j}-{1\over 3}\delta_{ij}\Phi^{,k}\Phi_{,k}\Big{)}\Big{(}1-{4\over c^{2}}\Phi\Big{)}.$ (60)
@@
$T^{0}_{0}=-\varrho c^{2}\Big{(}1+{1\over c^{2}}\Pi\Big{)},\quad T^{0}_{i}=0,\quad T^{i}_{0}=-{1\over ca}\varrho P^{i},\quad T^{i}_{j}=p\delta^{i}_{j}+\Big{(}1+{2\over c^{2}}\Psi\Big{)}\Pi^{i}_{j}.$ (61)
We note that setting $q_{i}=0$ is a frame choice always allowed for any energy-momentum component. Under this choice, the flux or momentum of the component is encoded in the fluid four-vector $u_{a}$. As the four-vector becomes normal with $n_{i}\equiv 0$, we have the physical momentum of the component vanish; this is reflected in $T^{0}_{i}=(\mu+p)u^{0}u_{i}=0$ to 1PN order. This implies that the $\tau$-fluid quantities are measured by an observer comoving with the $\tau$-fluid. This happened as we imposed $\sigma=0=\Sigma$ as the physical and gauge conditions.
@@
Equation (67) is valid to 0PN order whereas the other two equations are valid to 1PN order. To 0PN order, Eq. (66) gives Eq. (38) with $\sigma=0$. To 0PN order, Eq. (67) is the same as Eq. (66), with $\Psi=\Phi$.
The fluid quantities in Eqs. (66)-(68) are the sum of other fluids besides the $\tau$-fluid; we may include the $\tau$-field to the collective fluid quantities as $\mu=\mu_{m}+\mu_{\tau}$, $p=p_{m}+p_{\tau}$, and $\varrho v_{i}=\varrho^{m}v^{m}_{i}+\varrho^{\tau}v^{\tau}_{i}$. For the $\tau$-fluid we have conservation equations in (62)-(63) or the equation of motion in Eq. (64). The fluid conservation equations to 1PN order are in Eqs. (B23) and (B24) of Hwang-Noh-axion-2023 . These are the complete set of equations valid to 1PN order with $\Sigma=0$ as the temporal gauge condition; for other choice of gauge conditions, see Sec. 6 in Hwang-Noh-Puetzfeld-2008 .